# Assessing the impact of variance heterogeneity and misspecification in mixed-effects location-scale models

### v1

@@ 4.3 Practice 3: Ignoring non-linear time trend in location model
Another common type of misspecification in everyday practice is represented by the choice of a linear relationship between the outcome mean and the time variable when actually a non-linear trend is present. We, therefore, explore how non-linearity would affect the estimates of interest. To this end, we alter the data generation mechanism by applying a sinusoidal transformation on age as follows:
@@
while we keep a linear relationship with the scale parameter. In this context, we compare the properly specified models with the common misspecification ignoring the non-linearity (while the scale is correctly specified):
@@
Figure 5: Estimated age effect for location (left panel) and scale (right panel). For each model, coverage is reported as percentage next to the corresponding boxplot.
@@
In this case, the misspecification of the functional form of age leads to a biased estimate of the effect of the covariate on the location model. As the scale relies on age and not a functional transformation, the scale model remains non-impacted.
@@ 4.4 Practice 4: Misspecifying the random effect structure
Figure 6: Estimated age effect for location (left panel) and scale (right panel). For each model, coverage is reported as percentage next to the corresponding boxplot.
Figure 7: Estimated standard deviation of random intercepts for location (left panel) and scale (right panel). For each model, coverage is reported as percentage next to the corresponding boxplot.
@@
When a non-zero random slope is neglected, the estimates of the age effect (that corresponds to the fixed slope or population effect) are centred around the true value, but the coverage is much lower than the nominal level. It is worth mentioning that missing the random slope in the scale model does not influence the fixed effect in the location model, but in the opposite case, the age effect in the scale model shows a loss of coverage. When random intercepts and slopes are independent, neglecting the random slopes induces no changes in the standard deviations of the random intercepts.
@@ 4.5 Practice 5: Misspecifying random effect distributions
This setting resulted in no difference in the fixed effect estimates, our analysis therefore focuses on the random intercepts in Figure 8.
Figure 8: Estimated standard deviation of random intercepts for location (left panel) and scale (right panel). For each model, coverage is reported as percentage next to the corresponding boxplot.
@@
When the true random effect distribution is t-Student, the standard deviation of the Gaussian distribution will be estimated so that the Gaussian distribution matches the original distribution, as illustrated in Figure 9. This results in similar estimated distributions despite misestimated standard deviations. This result is consistent with the LMM literature (e.g.16).
@@ 5 Discussion
In this work, we have presented a set of simulations to study the impact of misspecification on mixed-effect location-scale models, with a focus on modelling heteroscedasticity as opposed to the assumption of constant residual variance made by standard linear mixed models.
Our simulations offer valuable insights into MELSM. First, scale misspecification increases the variance of the location estimates but does not introduce bias. In contrast, location misspecification biases scale estimates. Therefore, all conclusions about covariate effects in the scale model are to be referred to the specific set of covariates and functional forms used in the location. Further, there is an interplay between the random effects in the location model and the residual variance. Different models for residual variance influence the standard deviations of the location random effects. Finally, misspecification of the model structure, such as neglecting a relevant covariate, leads to larger bias than misspecification of the functional form, as including relevant covariates helps account for part of their effects, even if the functional form is not correct. These insights show the critical importance of modelling residual variance, inviting statisticians to model this quantity even under imperfect knowledge of its structure.
As a simulation, our findings have limitations. While we have used a general set of parameters to improve the generalisability of our conclusions, more extreme parameter settings could depart from the results shown here. Similarly, our simulations rely on two random effects and do not consider random effects for B-spline coefficients. Our conclusion on the misspecification of random effect distribution considers a symmetric alternative to the normal distribution and not more skewed distributions, which could impact this finding. Finally, our work focuses on the distribution shape and not on the prediction of individual random effects, which the linear mixed models literature suggests being more affected by the misspecification (see 7, 6, 19 for reviews on the topic).
This work opens avenues for evaluating misspecifications in other location-scale models — for instance, the addition of time-varying or non-normally distributed covariates. Beyond the presented setting, MELSMs are often used as part of more complex models. Particularly, the study of joint models with within-individual variability [1, 10] would provide insights into the relationship between the misspecification of the longitudinal model and the association with the hazard in time-to-event data analysis.
@@
Ethics and Consent to Participate declarations. Not applicable.
@@
Funding. This work was supported by the UK Medical Research Council grant “Looking beyond the mean: what within-person variability can tell us about dementia, cardiovascular disease and cystic fibrosis” (MR/V020595/1). JKB was supported through the United Kingdom Medical Research Council programme (grant MC_UU_00002/5) and theme (grant MC_UU_00040/02 – Precision Medicine) funding. MP’s research is supported by the Ulverscroft Vision Research Group (UCL).
@@ Appendix A Why do we expect a biased location model to have a larger scale error?
Let us simplify the problem by considering the following generative process:
$Y=\beta X+\epsilon$
with $\epsilon\sim\mathcal{N}(0,\omega)$.
A biased scale model can be expressed as:
$Y=(\beta+\delta)X+\epsilon^{\prime}$
@@
with $\epsilon^{\prime}\sim\mathcal{N}(0,\omega^{\prime})$ and $\delta$ the bias from the true coefficient.
Computing the difference between the generative process and the biased model results in:
$\epsilon^{\prime}=\epsilon-\delta X$
with $X\sim MVN(0,\Sigma)$. And therefore, one can express the variance relation as:
$\omega^{\prime}=\omega+\delta^{T}\Sigma\delta$
@@
Note that the exact quantification of these quantities are dependent upon the correlation structure between covariates. However, as $\Sigma$ is positive semi-definite:
$\omega^{\prime}\geq\omega$
This means that the estimated variance of a biased model is larger than the variance of the data.

### v2

@@ 4.3 Practice 3: Ignoring non-linear time trend in location model
Another common type of misspecification in everyday practice is the choice of a linear relationship between the outcome mean and the time variable when, in fact, a non-linear trend is present. We, therefore, explore how non-linearity would affect the estimates of interest. To this end, we alter the data generation mechanism by applying a sinusoidal transformation on age as follows:
@@
while we keep a linear relationship with the scale parameter. In this context, we compare the properly specified models with the common misspecification of ignoring the non-linearity (while the scale is correctly specified):
@@
Figure 5: Estimated age effect for location (left panel) and scale (right panel). For each model, coverage is reported as a percentage next to the corresponding boxplot.
@@
In this case, the misspecification of the functional form of age leads to a biased estimate of the effect of the covariate on the location model. As the scale relies on age and not a functional transformation, the nonlinearity in the location model has little impact on the scale submodel.
@@ 4.4 Practice 4: Misspecifying the random effect structure
Figure 6: Estimated age effect for location (left panel) and scale (right panel). For each model, coverage is reported as a percentage next to the corresponding boxplot.
Figure 7: Estimated standard deviation of random intercepts for location (left panel) and scale (right panel). For each model, coverage is reported as a percentage next to the corresponding boxplot.
@@
When a non-zero random slope is neglected, the estimates of the age effect (corresponding to the fixed slope or population effect) are centered around the true value, but the coverage is much lower than the nominal level. It is worth mentioning that missing the random slope in the scale model does not influence the fixed effect in the location model, but in the opposite case, the age effect in the scale model shows a loss of coverage. When random intercepts and slopes are independent, neglecting the random slopes induces no changes in the standard deviations of the random intercepts.
@@ 4.5 Practice 5: Misspecifying random effect distributions
This setting resulted in no difference in the fixed effect estimates; our analysis, therefore, focuses on the random intercepts in Figure 8.
Figure 8: Estimated standard deviation of random intercepts for location (left panel) and scale (right panel). For each model, coverage is reported as a percentage next to the corresponding boxplot.
@@
When the true random effect distribution is t-Student, the standard deviation of the Gaussian distribution will be estimated so that the Gaussian distribution matches the original distribution, as illustrated in Figure 9. This results in similar estimated distributions despite misestimated standard deviations. This result is consistent with the LMM literature in which misspecification of random effects distribution has a limited impact on model estimates [schielzeth2020robustness].
5 Case Study: Modelling heteroscedasticity in the PBC dataset
To further illustrate the impact of these different practices on a study’s conclusions, we propose an analysis of the PBC dataset from the R package JMBayes2. This example is intended as a didactic demonstration rather than a substantive epidemiological study.
Dataset.
The dataset results from a clinical trial at Mayo Clinic studying the impact of D-penicillamine on primary biliary cirrhosis (PBC), an autoimmune condition affecting the liver. It comprises 1,945 longitudinal observations from 312 patients diagnosed with PBC who received care at the Mayo Clinic. These data have been foundational in understanding the progression of this chronic autoimmune liver disease and associated short-term survival [murtaugh1994primary]. In our analysis, we model log-transformed serum bilirubin levels, a marker of condition severity and progression used for diagnosis and prognosis, as a function of the normalized age at baseline, sex, and time (in years since baseline).
Empirical setting.
Our goal is to assess how failing to account for heteroscedasticity can affect model estimates. To this end, we compare a traditional LMM and MELSM, which only differ in the way they model the scale. Note that for flexible modelling of the time effect, we use splines (with default parameters as per the package brms), denoted by $s(\cdot)$. For fitting these models, we use 4 chains with a burn-in of 1000 iterations for a total of 2000 iterations.
  • LMM assumes homoscedasticity.
            log(bilirubin) ~  age + sex + plat + alb + s(year) + (1|id)
            log(omega) ~ 1
  • MELSM no covariates models heteroscedasticity through random effects alone.
            log(bilirubin) ~ age + sex + plat + alb + s(year) + (1|id)
            log(omega) ~ (1|id)
  • MELSM no slope ignore both random slopes.
            log(bilirubin) ~ age + sex + plat + alb + s(year) + (1|id)
            log(omega) ~ age + sex + plat + alb + s(year) + (1|id)
  • MELSM no slope scale ignores the scale random slope.
            log(bilirubin) ~ age + sex + plat + alb + s(year) + (1 + age|id)
            log(omega) ~ age + sex + plat + alb + s(year) + (1|id)
  • MELSM slopes models the location and scale as follows.
            log(bilirubin) ~ age + sex + plat + alb + s(year) + (1 + age|id)
            log(omega) ~ age + sex + plat + alb + s(year) + (1 + age|id)
Results.
Figure 10 displays the fixed effects estimated by all the previous approaches. Critically, there are limited differences in the fixed effects estimated by the different alternatives of MELSM, echoing the finding that misspecification of the scale does not bias fixed effects as long as the heteroscedasticity is captured. While LMM presents similar estimates for the majority of the covariates, the estimated spline for years past since diagnosis differs, presenting a steeper impact. This example illustrates that ignoring error variance can impact inference and underscores the practical importance of explicitly modelling heteroscedasticity when it is present.
Figure 10: Estimated location fixed effects in the PBC dataset.
Figure 11 presents the scale fixed effects for the three models considering the different covariates in the scale model. The estimates are similar across models, indicating a limited impact of introducing random slopes in the specification of the scale submodel.
Figure 11: Estimated scale fixed effects in the PBC dataset.
Figure 12 presents the estimated standard deviation of the random effects for all MELSM models considering heteroscedasticity. The standard deviation of the scale random effects presents larger values for MELSM with no covariates, as this model does not control for any other covariates. Once covariates or random slopes are accounted for in the scale submodel, the standard deviation of the estimated random intercepts decreases. Importantly, note that the posterior distribution for the scale random intercept lies above 0, providing therefore evidence of heteroscedasticity.
Figure 12: Estimated standard deviation of random intercepts for location (left panel) and scale (right panel) in the PBC dataset.
@@ 6 Discussion
In this work, we present a set of simulations to study the impact of misspecification on mixed-effect location-scale models, with a focus on modelling heteroscedasticity as opposed to the assumption of constant residual variance made by standard linear mixed models.
Our simulations offer valuable insights into MELSM. First, scale misspecification increases the variance of the location estimates but does not introduce bias. In contrast, location misspecification biases scale estimates. Therefore, all conclusions about covariate effects in the scale model are to be referred to the specific set of covariates and functional forms used in the location. Further, there is an interplay between the random effects in the location model and the residual variance. Different models for residual variance influence the standard deviations of the location random effects. Finally, misspecification of the model structure, such as neglecting a relevant covariate, leads to larger bias than misspecification of the functional form, as including relevant covariates helps account for part of their effects, even if the functional form is not correct. These insights highlight the critical importance of modelling residual variance, inviting statisticians to model this quantity even under imperfect knowledge of its structure.
As a simulation, our findings have limitations. While we have used a general set of parameters to improve the generalisability of our conclusions, more extreme parameter settings could depart from the results shown here. Similarly, our simulations rely on two random effects and do not account for random effects in the B-spline coefficients. Our conclusion regarding the misspecification of the random effect distribution considers a symmetric alternative to the normal distribution, rather than more skewed distributions, which could impact this finding. Further, the proposed simulations assume an uninformative visit process, which in many real cases (e.g. electronic health records) might be unrealistic. Future developments in this area are needed to build joint models for the within-individual variability and the informative observation process, extending, for example, the methodologies illustrated in pullenayegum2016longitudinal and garrett2024biasmixedmodelsanalysing. Finally, our work focuses on the distribution shape, rather than the prediction of individual random effects, which the linear mixed models literature suggests is more affected by misspecification (see lebeau2018model, hui2021random, vu2025random for reviews on the topic).
Our findings may have implications for other longitudinal frameworks dependent on MELSM. Beyond the presented setting, MELSMs are often used as part of more complex models. Particularly, the study of joint models with within-individual variability [barrett2019estimating, palma2025bayesian] would provide insights into the relationship between the misspecification of the longitudinal model and the association with the hazard in time-to-event data analysis.
@@
This version of the article has been accepted for publication, after peer review, but is not the Version of Record and does not reflect post-acceptance improvements or any corrections. The Version of Record is available online at: http://dx.doi.org/10.1186/s12874-025-02755-3.
Ethics approval and consent to participate. Not applicable.
@@
Funding. This work was supported by the UK Medical Research Council grant “Looking beyond the mean: what within-person variability can tell us about dementia, cardiovascular disease and cystic fibrosis” (MR/V020595/1). JKB was supported through the United Kingdom Medical Research Council programme (grant MC_UU_00002/5) and theme (grant MC_UU_00040/02 – Precision Medicine) funding. MP’s research is currently supported by the Ulverscroft Vision Research Group (UCL).
@@
Acknowledgements. The authors would like to thank the reviewers for their valuable feedback throughout the review process.
@@ Appendix A Why do we expect a biased location model to have a larger scale error?
Let us simplify the problem by considering the following generative process: Y = βX + ϵ with $\epsilon\sim\mathcal{N}(0,\omega)$.
@@
A biased scale model can be expressed as: Y = (β+ δ) X + ϵ’ with $\epsilon^{\prime}\sim\mathcal{N}(0,\omega^{\prime})$ and $\delta$ the bias from the true coefficient.
@@
Computing the difference between the generative process and the biased model results in: ϵ’ = ϵ- δX with $X\sim MVN(0,\Sigma)$. And therefore, one can express the variance relation as: ω’ = ω+ δ^T Σδ Note that the exact quantification of these quantities is dependent upon the correlation structure between covariates. However, as $\Sigma$ is positive semi-definite: ω’ ≥ω This means that the estimated variance of a biased model is larger than the variance of the data.
Appendix B Additional metrics
In addition to the coverage and estimates presented in the main text, we present the additional metrics of the average credible interval width ($\bar{CI}(\theta)$), the standard deviation of the estimates $SD(\theta)$, and average posterior standard deviations $\bar{\epsilon}(\theta)$ for the same parameters of interest across the different replications.
B.1 Practice 1: Ignoring heteroscedasticity
These tables echo the same conclusions as the main text, with the correctly specified model presenting smaller credible interval widths, deviations of the estimate, and average standard errors by an order of magnitude. Increasing numbers of individuals or points also result in improvements across all metrics.
SDSD($\beta^{y}[Age]$) $\bar{CI}$($\beta^{y}[Age]$) $\bar{\epsilon}$($\beta^{y}[Age]$)
Models
100 Correct 0.017 0.063 0.016
300 Correct 0.009 0.035 0.009
500 Correct 0.007 0.026 0.007
1000 Correct 0.005 0.018 0.005
Table 3: Measures of errors of age effect on location for increasing number of individuals.
SDSD($\beta^{\omega}[Age]$) $\bar{CI}$($\beta^{\omega}[Age]$) $\bar{\epsilon}$($\beta^{\omega}[Age]$)
Models
100 Correct 0.024 0.092 0.023
300 Correct 0.013 0.053 0.013
500 Correct 0.010 0.041 0.010
1000 Correct 0.008 0.029 0.007
Table 4: Measures of errors of age effect on scale for increasing number of individuals.
SDSD($\beta^{y}[Age]$) $\bar{CI}$($\beta^{y}[Age]$) $\bar{\epsilon}$($\beta^{y}[Age]$)
Models
5 Correct 0.021 0.082 0.021
10 Correct 0.013 0.055 0.014
20 Correct 0.009 0.037 0.009
Table 5: Measures of errors of age effect on location for increasing number of points.
SDSD($\beta^{\omega}[Age]$) $\bar{CI}$($\beta^{\omega}[Age]$) $\bar{\epsilon}$($\beta^{\omega}[Age]$)
Models
5 Correct 0.039 0.140 0.036
10 Correct 0.022 0.084 0.021
20 Correct 0.014 0.055 0.014
Table 6: Measures of errors of age effect on scale for increasing number of points.
B.2 Practice 2: Misspecifying location and scale in MELSM
Similarly, this section illustrates how different misspecifications affect the different fixed effects with regard to these metrics. Interestingly, the model that considers all covariates and the correctly specified one are indistinguishable from these metrics.
SDSD($\beta^{y}[Age]$) $\bar{CI}$($\beta^{y}[Age]$) $\bar{\epsilon}$($\beta^{y}[Age]$)
Models
Correct 0.011 0.043 0.011
All 0.011 0.043 0.011
Mis. $\omega$ 0.018 0.056 0.014
No $u^{\omega}$ 0.017 0.065 0.017
No $u^{\omega}$ + Mis. $\omega$ 0.044 0.109 0.028
Table 7: Measures of errors of age effect on location.
SDSD($\beta^{\omega}[Age]$) $\bar{CI}$($\beta^{\omega}[Age]$) $\bar{\epsilon}$($\beta^{\omega}[Age]$)
Models
Correct 0.016 0.065 0.017
All 0.016 0.065 0.017
Mis. $y$ 0.017 0.068 0.017
No $u^{\omega}$ 0.032 0.055 0.014
Table 8: Measures of errors of age effect on scale.
SDSD($\sigma_{y}$) $\bar{CI}$($\sigma_{y}$) $\bar{\epsilon}$($\sigma_{y}$)
Models
Correct 0.054 0.217 0.055
All 0.054 0.218 0.056
Mis. $y$ 0.054 0.218 0.056
Mis. $\omega$ 0.058 0.234 0.060
No $u^{\omega}$ 0.055 0.219 0.056
No $u^{\omega}$ + Mis. $\omega$ 0.071 0.266 0.068
Table 9: Measures of errors of standard deviation of random intercepts for location.
SDSD($\sigma_{\omega}$) $\bar{CI}$($\sigma_{\omega}$) $\bar{\epsilon}$($\sigma_{\omega}$)
Models
Correct 0.031 0.123 0.031
All 0.031 0.124 0.032
Mis. $y$ 0.029 0.109 0.028
Mis. $\omega$ 0.041 0.185 0.047
Table 10: Measures of errors of standard deviation of random intercepts for scale.
B.3 Practice 3: Ignoring non-linear time trend in location model
The following tables illustrate these metrics when the correctly specified model depends on the sine of age. Interestingly, these metrics do not capture the bias in the estimated fixed effect, even if the improperly specified model shows narrower parameter estimates. This aligns with the intuition that more complex models may lead to larger standard deviations for the different parameters of interest.
SDSD($\beta^{y}[Age]$) $\bar{CI}$($\beta^{y}[Age]$) $\bar{\epsilon}$($\beta^{y}[Age]$)
Models
Correct 0.015 0.058 0.015
Non sinus 0.014 0.047 0.012
Table 11: Measures of errors of age effect on location.
SDSD($\beta^{\omega}[Age]$) $\bar{CI}$($\beta^{\omega}[Age]$) $\bar{\epsilon}$($\beta^{\omega}[Age]$)
Models
Correct 0.016 0.065 0.017
Non sinus 0.016 0.065 0.017
Table 12: Measures of errors of age effect on scale.
B.4 Practice 4: Misspecifying the random effect structure
In this setting, the metrics again do not capture the estimate bias, as the correctly specified model presents worse performance on the random effect estimates despite presenting less bias.
The tables show the effect of removing the random slope for age in each of the two submodels. We observe that, once we remove the random slope coefficient from the location submodel, the corresponding fixed effect is affected by low coverage and the estimates of the metrics in the tables are different from those in the correct model. A similar behaviour is observed for the random slope in the scale submodel.
The metrics for the standard deviation of the random intercept in the location and scale submodels remains essentially unaltered after removing the random slopes.
SDSD($\beta^{y}[Age]$) $\bar{CI}$($\beta^{y}[Age]$) $\bar{\epsilon}$($\beta^{y}[Age]$)
Models
Correct 0.031 0.120 0.031
No $u^{\omega}_{age}$ 0.032 0.121 0.031
No slopes 0.046 0.066 0.017
Table 13: Measures of errors of age effect on location.
SDSD($\beta^{\omega}[Age]$) $\bar{CI}$($\beta^{\omega}[Age]$) $\bar{\epsilon}$($\beta^{\omega}[Age]$)
Models
Correct 0.029 0.114 0.029
No $u^{\omega}_{age}$ 0.033 0.068 0.017
No slopes 0.032 0.068 0.017
Table 14: Measures of errors of age effect on scale.
SDSD($\sigma_{y}$) $\bar{CI}$($\sigma_{y}$) $\bar{\epsilon}$($\sigma_{y}$)
Models
Correct 0.060 0.236 0.060
No $u^{\omega}_{age}$ 0.060 0.237 0.061
No slopes 0.053 0.223 0.057
Table 15: Measures of errors of standard deviation of random intercepts for location.
SDSD($\sigma_{\omega}$) $\bar{CI}$($\sigma_{\omega}$) $\bar{\epsilon}$($\sigma_{\omega}$)
Models
Correct 0.034 0.132 0.034
No $u^{\omega}_{age}$ 0.034 0.129 0.033
No slopes 0.032 0.125 0.032
Table 16: Measures of errors of standard deviation of random intercepts for scale.
B.5 Practice 5: Misspecifying random effect distributions
In this final experiment, the proposed metrics reveal limited differences between the two distribution specifications of the random effects (except for $SD(\sigma_{y})$, which is dependent on the different distribution parameterisation). This is consistent with the findings on the limited impact of distribution misspecification.
SDSD($\sigma_{y}$) $\bar{CI}$($\sigma_{y}$) $\bar{\epsilon}$($\sigma_{y}$)
Models
Correct 0.147 0.345 0.093
Gaussian 0.190 0.329 0.086
Table 17: Measures of errors of standard deviation of random intercepts for location.
SDSD($\sigma_{\omega}$) $\bar{CI}$($\sigma_{\omega}$) $\bar{\epsilon}$($\sigma_{\omega}$)
Models
Correct 0.068 0.190 0.051
Gaussian 0.142 0.182 0.047
Table 18: Measures of errors of standard deviation of random intercepts for scale.